# TrialSynth: Generation of Synthetic Sequential Clinical Trial Data

### v1

@@ Abstract
Analyzing data from past clinical trials is part of the ongoing effort to optimize the design, implementation, and execution of new clinical trials and more efficiently bring life-saving interventions to market. While there have been recent advances in the generation of static context synthetic clinical trial data, due to both limited patient availability and constraints imposed by patient privacy needs, the generation of fine-grained synthetic time-sequential clinical trial data has been challenging. Given that patient trajectories over an entire clinical trial are of high importance for optimizing trial design and efforts to prevent harmful adverse events, there is a significant need for the generation of high-fidelity time-sequence clinical trial data. Here we introduce TrialSynth, a Variational Autoencoder (VAE) designed to address the specific challenges of generating synthetic time-sequence clinical trial data. Distinct from related clinical data VAE methods, the core of our method leverages Hawkes Processes (HP), which are particularly well-suited for modeling event-type and time gap prediction needed to capture the structure of sequential clinical trial data. Our experiments demonstrate that TrialSynth surpasses the performance of other comparable methods that can generate sequential clinical trial data, in terms of both fidelity and in enabling the generation of highly accurate event sequences across multiple real-world sequential event datasets with small patient source populations when using minimal external information. Notably, our empirical findings highlight that TrialSynth not only outperforms existing clinical sequence-generating methods but also produces data with superior utility while empirically preserving patient privacy.
@@ 1 Introduction
The data generated from past clinical trials represent a valuable resource for informing drug development Fu et al. (2022); Chen et al. (2024b); Wang et al. (2023b); Chen et al. (2024c) and increasing the speed at which vital life-saving drugs arrive to market Friedman et al. (2015); Downing et al. (2014) While the potential value of clinical trial data is high, these data are often inaccessible due to patient privacy concerns and legal constraints (Shafquat et al., 2023; Lu, 2018; Hu & Li, 2024). The generation of high-quality synthetic clinical trial data that captures the properties of real data while simultaneously protecting patient privacy is increasingly being seen as a strategy for sharing and applying these data in drug development applications James et al. (2021).
Though proposed methods for generating synthetic clinical trial data have focused on static context information for each subject (e.g., demographics) Hernandez et al. (2022), many of the highest value applications, including control arm augmentation Thorlund et al. (2020) require generating synthetic time-sequential event data that has high fidelity (Beigi et al., 2022). However, developing a high-quality model for sequential trial data can be more complicated than data-rich tasks in computer vision or natural language processing, due to the small sample size of training datasets available, which is less common in other applications of generative models.
To address these challenges, we propose TrialSynth. This method makes use of Hawkes processes, which are statistical models that are specialized for event-type and time gap prediction (Hawkes, 1971; Zuo et al., 2020), as well as Variational Autoencoders (VAEs) (Kingma & Welling, 2013), a proven generative framework that has worked well for static clinical trial data synthesis Das et al. (2023). We empirically demonstrate that combining these two classical approaches leads to an algorithm that is capable of generating sequential event synthetic data even on small amounts of clinical trial data.
@@
    We introduce TrialSynth–a model that combines Variational Autoencoder + Hawkes Process that is both able to generate sequential event clinical trial data and supports a high level of control, allowing users to specify specific event types to generate.
@@ 2 Related Work
Synthetic Data Generation as a research area has been quickly garnering attention from the research community, with examples such as CTGAN (Xu et al., 2019), CTabGan (Zhao et al., 2022), TabDDPM (Kotelnikov et al., 2023), TWIN-GPT (Wang et al., 2024), the Synthetic Data Vault [^1] https://docs.sdv.dev/sdv/ (Patki et al., 2016), and more. However, most of these models, such as TabDDPM and CTGAN, are focused on explicitly generating tabular data with no time component; or, in the case of SDV’s ParSynthesizer (Zhang et al., 2022), it is relatively simple and may be approximated with a GRU or LSTM model.
@@
Trial Patient Generation is a research area that has become popular. In Electronic Healthcare Record (EHR) generation (Choi et al., 2017; Chen et al., 2021; Wu et al., 2022; Das et al., 2023; Wang et al., 2023b; Liu et al., 2023; Theodorou et al., 2023; Wang et al., 2024), the model usually only focuses on generating the order at which certain clinical events happen (i.e., the diagnosis code of next patient visit), as opposed to generating the specific times of the visits as well. For example, Das et al. (2023); Wang et al. (2024) generates a digital twin of an input patient event sequence via a VAE and a cross-modality model, but cannot handle event timestamp generation. TrialSynth extends this line of previous work to include the specific timestamps on which these events occur, as well as the order. Theodorou et al. (2023) created a strong patient EHR generation baseline, but relies on a high amount of training data (929,268 and 46,520 patients in outpatient and inpatient datasets respectively). However, in a single clinical trial, all of our datasets contain less than 1000 patients, which makes HALO difficult to run. TrialSynth is designed for and performs well on small clinical trial datasets, particularly if the event types are known.
Hawkes Processes combined with VAEs is an area of research that is particularly appealing for our scenario. Traditional Hawkes Processes are a double stochastic point process. It assumes that past events can temporarily raise (never lower) the probability of future events, assuming that such excitation is positive, additive over past events, and exponentially decaying with time: $\lambda(t):=\left(\mu+\sum_{\{t_{j},t_{j}<t\}}e^{-\delta(t-t_{j})}\right)$ where $\delta$ is a decaying rate, the occurrence of a single event occurring at $t$ has intensity $\lambda(t)$, calculated as the sum of base intensity $\mu$ as well as all previously occurring events $t_{j}<t$.
@@
This inherent capability of modeling events and their time occurrences makes Hawkes Processes highly suitable for event prediction. Previous work explores variational Hawkes processes in the context of event prediction for (disease progression (Chen et al., 2024a) and social events sequences Pan et al. (2020), but they rely on the context of previous ground truth observations as well as the hidden state. Another work (Lin et al., 2021) explores using variational approaches to disentangle multivariate Hawkes Process for event type prediction, but it also relies on knowing the ground truth to predict the next timestep. This limitation is a major roadblock in a full synthetic data generation setting. Because of this, there is leaking of information from ground truth event occurrences. This information leakage is not permitted in our task, which is a fully generative setting from the embedding space. Additionally, we also employ a more recent version of the neural Hawkes Process, namely, the Transformer Hawkes Process Zuo et al. (2020). To the best of our knowledge, TrialSynth is the first to extend Hawkes models to full patient event generation from a single embedding.
@@ 3 TrialSynth
Figure 2: Diagram of the TrialSynth Encoder-Decoder structure. Here, the model input is the real patient event sequence + time, which is used to train a VAE model to the same output time + event sequence. The event sequence length for each event (if not known) is also predicted. Optional inputs (in dashed lines) include the ground-truth knowledge of which event types to generate and/or event lengths. The input observations are encoded first by a specific event encoder, and then embeddings are mapped to the patient embedding space via an MLP. The opposite process occurs for decoding from the patient embedding.
@@ 3.1 Encoding and Decoding Hawkes Processes
Neural Hawkes Process (Mei & Eisner, 2017) was proposed to generalize the traditional Hawkes Process. Let us describe the $\lambda(t)$, the intensity function of any event occurring at time $t$.
@@
$\lambda_{k}(t)$ is the intensity function for the event $k\in\mathcal{K}$ occurring, $K=|\mathcal{K}|$ is the total number of event types, $h(t)$ are the hidden states of the event sequence obtained by a Transformer encoder, and $\bm{W}_{k}^{\top}$ are learned weights that calculate the significance of each event type at time $t$. $f_{k}(c)=\beta_{k}\log(1+e^{\frac{x}{\beta_{k}}})$ is the softplus function with parameter $\beta_{k}$. The output of $f_{k}(x)$ is always positive. Note that the positive intensity does not mean that the influence is always positive, as the influence of previous events are calculated through $\bm{W}_{k}^{\top}\bm{h}(t)$. If there is an event occurring at time $t$, then the probability of event $k$ is $P(k_{t}=k)=\frac{\lambda_{k}(t)}{\lambda(t)}$. Furthermore, the log-likelihood is:
@@
The encoder model $E_{\texttt{TrialSynth}}(\mathcal{H}_{i})\rightarrow\hat{\bm{\mu}},\hat{\bm{\sigma}}$ takes in the original event types and times, and predicts the mean and standard deviation to sample hidden state vector $\bm{z}\sim Normal(\hat{\bm{\mu}},\hat{\bm{\sigma}})$. This is trained to be close to the $Normal(\bm{0},\bm{1})$.
However, the decoder model is more complicated. We train the decoder to maximize the likelihood of the input Hawkes Process. I.e. the input is the ground truth event type and time-step sequence, and the autoencoder reconstructs it from $\bm{z}$. For our purposes, we adapt this into an autoregressive decoding scheme similar to the language modeling task.
@@
is a hidden vector $\bm{z}$ and a sequence of ground truth event types and event times. It is tasked with predicting the next type of event $\hat{k}$, the next event time $\hat{t}$, and the intensity function $\lambda$ that measures the probability of an event occurring. $\lambda$ is necessary to compute the likelihood $P_{\theta}((t_{1},k_{1}),\dots,(t_{i},k_{i}),|\bm{z})$.  [^2] Things are abstracted out for clarity, please see Appendix for details. Furthermore, we follow Transformer Hawkes Process’s approach of also adding mean squared error losses to the time: $time\_loss=\|t-\hat{t}\|^{2}$ and cross-entropy loss of the predicted $type\_loss=-\sum_{c=1}^{|\mathcal{K}|}k\log(p_{k})$
At inference time, the input to the decoder is only $z$, and we autoregressively decode the predicted event types and times. To predict next time and event tuple $(\hat{t}_{i},\hat{k}_{i})$, the input is the previously predicted times and events $\{(\hat{t}_{1},\hat{k}_{1}),\dots,(\hat{t}_{i-1},\hat{k}_{i-1}))\}$). (each predicted time and event is repeatedly appended to the input).
We find that greedy auto-regressive sampling $\hat{k}=\operatorname*{arg\,max}(\bm{\hat{k}})$ of the next event type does not result in good sequences, and probabilistic sampling based on the raw logit score will result in higher utility sequences. Finally, we note that we can control for the generation of events that are similar to the original patient by first encoding the original patient and then sampling around it, a benefit of the probabilistic nature of the VAE latent space $z$. Otherwise, it would be impossible to correspond the original labels to the synthetic data.
@@
Table 1: A description of all the real-world datasets used in the evaluation. All trial data was obtained from Project Data Sphere (Green et al., 2015). Num Rows refers to the raw number of data points in the trial. Num Subj refers to the total number of patients. Num Events denotes the total number of unique events. Events / Subj denotes the average number of events that a patient experiences. Positive Label Proportion denotes the percentage of patients that did not experience the death event.
@@ 3.2 Event Type Information
We also propose 2 variants of TrialSynth. In some applications, such as clinical trial patient modeling (Wang & Sun, 2022a; Fu et al., 2022; Chen et al., 2024b, c; Chang et al., 2019; Das et al., 2023), we may be interested in an event sequence with known event types, that is, the model only needs to generate the timestamps at which events occur. This is to address the concern of subject fidelity, that is, the generated subject must be significantly similar to the original subject in order for the generated data to be useful; therefore, knowing which events occur in a subject to generate a similar subject would not be unreasonable. Along with the “Events Unknown” model that has no assumptions, we also propose the “Events Known” model was created to enforce ONLY simulating specific events, without considering all events (which may be irrelevant and confuse the generator).
To accommodate this framework, we train $num\_event$ Transformer Hawkes Process expert encoders and decoders that only model a single event. Since the event type is known, we can index in the specific encoder/decoder pair as given by the event index (shown in Figure 2). Finally, all independent event time predictions over all known types are combined via their predicted event times to create a final multi-event sequence prediction. Since the VAE model requires a single embedding to sample a latent vector, we sum the patient embedding output of all expert event encoders and pass this joint embedding to the expert decoders. The decoder is trained to generate the predicted time and type sequence of its respective expert event.
Note that we have the sum of the expert decoders’ latent patient embeddings due to the limitation of having a fully generative model. It would be unfair to have latent spaces of varying sizes for each patient.
@@ 3.3 Sequence Length Prediction
Given known subset of events $\mathcal{K}^{\prime}\in\mathcal{K}$, we generate event sequences such that $S=\{(t_{j},k_{j});j=1,\dots,L;k_{j}\in\mathcal{K}^{\prime}\},$ where $L$ is learned from the data. Note that by default, all variants of TrialSynth autoregressively continue to predict event types and times until it reaches its own predicted stopping length. Like other synthesizers (Zhang et al., 2022), we also experiment with using a max-length criterion, such that the event generation still stops at a certain $L$. We explore 2 cases. (1) If the event type is not known, we simply generate sequence $S=\{(t_{j},k_{j});j=1,\dots,L^{\prime}\}$ where $L^{\prime}$ is a prespecified sequence length. (2) If the event type is known, then we allow the model to know the specific max lengths $L_{k}^{\prime}$ for each specific event type. Let $L^{\prime}=\sum_{k_{i}\in\mathcal{K}^{\prime}}L_{k_{i}}^{\prime}$ We then generate a sequence $S=\{(t_{j},k_{j});j=1,\dots,L^{\prime};\left(\sum_{j=1}^{L^{\prime}}\mathbbm{1}(k_{j}=k_{i})\right)=L_{k_{i}}^{\prime};\forall k_{i}\in\mathcal{K}^{\prime}\}$

### v2

@@ Abstract
Analyzing data from past clinical trials is part of the ongoing effort to optimize the design, implementation, and execution of new clinical trials and more efficiently bring life-saving interventions to market. While there have been recent advances in the generation of static context synthetic clinical trial data, due to both limited patient availability and constraints imposed by patient privacy needs, the generation of fine-grained synthetic time-sequential clinical trial data has been challenging. Given that patient trajectories over an entire clinical trial are of high importance for optimizing trial design and efforts to prevent harmful adverse events, there is a significant need for the generation of high-fidelity time-sequence clinical trial data. Here we introduce TrialSynth, a Variational Autoencoder (VAE) designed to address the specific challenges of generating synthetic time-sequence clinical trial data. Distinct from related clinical data VAE methods, the core of our method leverages Hawkes Processes (HP), which are particularly well-suited for modeling event-type and time gap prediction needed to capture the structure of sequential clinical trial data. Our experiments demonstrate that TrialSynth surpasses the performance of other comparable methods that can generate sequential clinical trial data at varying levels of fidelity / privacy tradeoff, enabling the generation of highly accurate event sequences across multiple real-world sequential event datasets with small patient source populations. Notably, our empirical findings highlight that TrialSynth not only outperforms existing clinical sequence-generating methods but also produces data with superior utility while empirically preserving patient privacy.
@@ 1 Introduction
The data generated from past clinical trials represent a valuable resource for informing drug development [17, 8, 42, 9] and increasing the speed at which vital life-saving drugs arrive to market [15, 13] While the potential value of clinical trial data is high, these data are often inaccessible due to patient privacy concerns and legal constraints [35, 27, 22]. The generation of high-quality synthetic clinical trial data that captures the properties of real data while simultaneously protecting patient privacy is increasingly being seen as a strategy for sharing and applying these data in drug development applications [23].
Though proposed methods for generating synthetic clinical trial data have focused on static context information for each subject (e.g., demographics) [21], many of the highest value applications, including control arm augmentation [37] require generating synthetic time-sequential event data that has high fidelity [4, 46]. However, developing a high-quality model for sequential trial data can be more complicated than data-rich tasks in computer vision or natural language processing, due to the small sample size of training datasets available, which is less common in other applications of generative models.
To address these challenges, we propose TrialSynth. This method makes use of Hawkes processes, which are statistical models that are specialized for event-type and time gap prediction [20, 51], as well as Variational Autoencoders (VAEs) [24], a proven generative framework that has worked well for static clinical trial data synthesis [12]. We empirically demonstrate that combining these two classical approaches leads to an algorithm that is capable of generating sequential event synthetic data even on small amounts of clinical trial data.
@@
    We introduce TrialSynth–a model that combines Variational Autoencoder + Hawkes Process that is both able to generate sequential event clinical trial data and supports a high level of control, allowing users to specify specific event types and variance levels to generate (https://github.com/chufangao/TrialSynth).
@@ 2 Related Work
Synthetic Data Generation as a research area has been quickly garnering attention from the research community, with examples such as CTGAN [45], CTabGan [50], TabDDPM [25], TWIN-GPT [38], the Synthetic Data Vault [^1] https://docs.sdv.dev/sdv/ [33], and more. However, most of these models, such as TabDDPM and CTGAN, are focused on explicitly generating tabular data with no time component; or, in the case of SDV’s ParSynthesizer [48], it is relatively simple and may be approximated with a GRU or LSTM model.
@@
Trial Patient Generation is a research area that has become popular. In Electronic Healthcare Record (EHR) generation [10, 7, 43, 12, 42, 28, 36, 38], the model usually only focuses on generating the order at which certain clinical events happen (i.e., the diagnosis code of next patient visit), as opposed to generating the specific times of the visits as well. For example, [12, 38] generates a digital twin of an input patient event sequence via a VAE and a cross-modality model, but cannot handle event timestamp generation. TrialSynth extends this line of previous work to include the specific timestamps on which these events occur, as well as the order. [36] created a strong patient EHR generation baseline, but relies on a high amount of training data (929,268 and 46,520 patients in outpatient and inpatient datasets respectively). However, in a single clinical trial, all of our datasets contain less than 1000 patients, which makes HALO difficult to run. TrialSynth is designed for and performs well on small clinical trial datasets, particularly if the event types are known.
Hawkes Processes combined with VAEs is an area of research that is particularly appealing for our scenario. We employ the Transformer Hawkes Process [51] for our data generation modeling. To the best of our knowledge, TrialSynth is the first to extend Hawkes models to full patient event generation from a single embedding. Unlike the Hawkes Process, it relaxes the assumption that past events can never lower the probability of future events, and performs much better on real world data.
@@
This inherent capability of modeling events and their time occurrences makes Hawkes Processes highly suitable for event prediction. Previous work explores variational Hawkes processes in the context of event prediction for (disease progression [6] and social events sequences [32], but they rely on the context of previous ground truth observations as well as the hidden state. Another work [26] explores using variational approaches to disentangle multivariate Hawkes Process for event type prediction, but it also relies on knowing the ground truth to predict the next timestep. This limitation is a major roadblock in a full synthetic data generation setting. Because of this, there is leaking of information from ground truth event occurrences. This information leakage is not permitted in our task, which is a fully generative setting from the embedding space.
@@ 3 TrialSynth
Figure 2: Diagram of the TrialSynth Encoder-Decoder structure. Here, the model input is the real patient event sequence + time, which trains a VAE model to the same output time + event sequence. The event sequence length for each event is also predicted. The transformer encoder processes each input timestep, then output embeddings are individually transformed to the z-latent space via a neural network. Sampling and decoding occur from each timestep-specific z-latent representation.
@@ 3.1 Encoding and Decoding Hawkes Processes
Neural Hawkes Process [30] was proposed to generalize the traditional Hawkes Process. Let us describe the $\lambda(t)$, the intensity function of any event occurring at time $t$.
@@
Encoder: The encoder model $E_{\texttt{TrialSynth}}(\mathcal{H}_{i})\rightarrow\hat{\bm{\mu}},\hat{\bm{\sigma}}$ takes in the original event types and times, and predicts the mean and standard deviation to sample hidden state vector at time $\bm{z_{t}}$ at each timestep $t$. These $\bm{z_{t}}$ are concatenated to form $\bm{z}\sim Normal(\hat{\bm{\mu}},\hat{\bm{\sigma}})$. $\bm{z}$ is trained to be close to the $Normal(\bm{0},\bm{1})$ via ELBO.
Decoder: We train the decoder to maximize the likelihood of the input Hawkes Process. I.e. the input is the ground truth event type and time-step sequence, and the autoencoder reconstructs it from $\bm{z}$. For our purposes, we adapt a decoding scheme similar to HALO [36].
@@
is a hidden vector $\bm{z}$ and a sequence of ground truth event types and event times. It is tasked with predicting the next type of event $\hat{k}$, the next event time $\hat{t}$, and the intensity function $\lambda$ that measures the probability of an event occurring. $\lambda$ is necessary to compute the likelihood $P_{\theta}((t_{1},k_{1}),\dots,(t_{i},k_{i}),|\bm{z})$.  [^2] Please see Appendix for details. Furthermore, we follow Transformer Hawkes Process’s approach of also adding mean squared error losses to the time: $time\_loss=\|t-\hat{t}\|^{2}$ and cross-entropy loss of the predicted $type\_loss=-\sum_{c=1}^{|\mathcal{K}|}k\log(p_{k})$
At inference time, the input to the decoder is only $z$, and we decode the predicted event types and times. To predict next time and event tuple $(\hat{t}_{i},\hat{k}_{i})$, the input is the previously predicted times and events $\{(\hat{t}_{1},\hat{k}_{1}),\dots,(\hat{t}_{i-1},\hat{k}_{i-1}))\}$). (each predicted time and event is repeatedly appended to the input).
Finally, we note that we can control for the generation of events that are similar to the original patient by first encoding the original patient and then sampling around it, a benefit of the probabilistic nature of the VAE latent space $z$. Otherwise, it would be impossible to correspond the original labels to the synthetic data. For all experiments in this work, we take a random sample of the latent vector $\bm{z}$ to reconstruct our patient. Otherwise, our task would collapse down to a straightforward autoencoder task.
3.2 Final Loss Terms
Finally, we write the final loss as
$L=L_{hawkes}+L_{elbo}+L_{length}$
The $L_{hawkes}$ is the log-likelihood of the sequence given the Hawkes process above. The $L_{elbo}$ is the VAE loss of the hidden vector KL divergence from a standard Gaussian, the mean-squared error reconstruction loss of the event times, and the cross-entropy loss of the event types. Finally, we additionally add $L_{length}$ to ensure the model learns proper sequence lengths (described in section 3.4).
@@ 3.2 Final Loss Terms
Table 1: A description of all the real-world datasets used in the evaluation. All trial data was obtained from Project Data Sphere [19]. Num Rows refers to the raw number of data points in the trial. Num Subj refers to the total number of patients. Num Events denotes the total number of unique events. Events / Subj denotes the average number of events that a patient experiences. Positive Label Proportion denotes the percentage of patients that did not experience the death event.
@@ 3.3 Event Type Information
We also propose 2 variants of TrialSynth. In some applications, such as clinical trial patient modeling [40, 17, 8, 9, 5, 12], we may be interested in an event sequence with known event types, that is, the model only needs to generate the timestamps at which events occur. This is to address the concern of subject fidelity, that is, the generated subject must be significantly similar to the original subject in order for the generated data to be useful; therefore, knowing which events occur in a subject to generate a similar subject would not be unreasonable. Along with the “Events Unknown” model that has no assumptions, we also propose the “Events Known” model was created to enforce ONLY simulating specific events, without considering all events (which may be too numerous and irrelevant to the current patient).
To accommodate TrialSynth (Events Known), we use the exact same model as TrialSynth (Events Unknown), but restrict the event type prediction module to only valid patient input event types at inference time. We retain the same training process for both models, since we do not want to restrict learning event type information at training.
@@ 3.4 Sequence Length Prediction
We generate event sequences $\{(t_{j},k_{j});j=1,\dots,L;k_{j}\in\mathcal{K}^{\prime}\},$ where length $L$ is also generated by TrialSynth. Taking inspiration from HALO [36], our generation process automatically appends an [END] event at the end of each of the patient events. Furthermore, in addition to the event loss from before, we add a cross-entropy loss term on specifically the [END] event.
@@ 4 Experiments
4.1 Datasets
We evaluated our models on 7 real-world clinical trial outcome datasets obtained from Project Data Sphere [^3] https://data.projectdatasphere.org/projectdatasphere/html/access [19, 16, 5, 8, 9]. Specifically, we chose the trials as outlined in Table 1. These datasets have shown to be effective evaluation datasets for tabular prediction [41, 39] and digital twin generation [12, 38]. Specifically, we use LC1 [31], BC1 [3], CC [2], BC2 [14], BC3 [44], VTE [1], LC2 [11, 18, 47]. A full description of the data is shown in Table 1. Each dataset contains events and the times at which they occur, e.g., medications and procedures, as well as some adverse events like vomiting etc. We use these datasets to predict if the subject experiences the death event, which is an external label. Note that TrialSynth does not require a fixed patient event sequence length.
4.2 Baseline Methods
One surprising challenge we found was that existing EHR methods and synthetic patient generation methods are not applicable to our specific task and dataset due to dataset size and lack of support for timestamp generation; therefore, we primarily compare against general sequential data generation methods.
We compared the following 7 models: First, the LSTM VAE is the same as our proposed model, except with an LSTM instead of a Transformer encoder. PARSynthesizer from is SDV, based on a conditional probabilistic auto-regressive (CPAR) model, is specifically tailored for synthesizing sequential event data and stands out due to its unique focus and accessible codebase. TabDDPM is a state-of-the-art tabular synthesizer using diffusion models, enhanced by adding time as a numerical column for our purposes. Despite not being explicitly designed for sequential data, it surpasses previous models like CTGAN in synthetic tabular data generation. Lastly, HALO, a hierarchical autoregressive language model, excels in synthesizing Electronic Health Records (EHR) but struggles with clinical trial datasets due to the limited size of the training data, highlighting the challenges in this domain.
TrialSynth (Events Unknown) is the VAE + Multivariate Hawkes Process that is trained without any assumptions. At training time, the task is to predict a patient’s events and timesteps given the latent vector. TrialSynth (Events Known) assumes that one knows which specific events occur for the Hawkes Model. This essentially just restricts the number of valid events in the prediction phase by patient’s unique events.
4.3 Utility Evaluation